# SeCap: Self-Calibrating and Adaptive Prompts for Cross-view Person Re-Identification in Aerial-Ground Networks

### v1

@@ 1 Introduction
Person re-identification (ReID), as the cornerstone of intelligent surveillance systems, fundamentally relies on accurately identifying individuals across different camera viewpoints and complex environmental changes [11, 12, 35, 33, 32]. However, traditional ReID methods are often limited to the same view, such as ground view [30, 39, 24, 31, 34] or aerial view [37, 14, 4, 10, 25], and fail to adequately address the challenges posed by extreme visual transformations and the integration of complementary information in cross-view scenarios (e.g., combining ground and aerial cameras). These challenges are more prevalent in real-world applications, thereby introducing the problem of the Aerial-Ground Person Re-Identification (AGPReID).
@@ 2.3 Aerial-Ground ReID Datasets
In the AGPReID field, the scarcity and limitations of datasets become the key factors constraining research progress. Compared to mature view-invariant datasets [40, 41], AGPReID datasets are scarce in number and far from meeting research needs in terms of real-world data scale. For instance, the number of identity IDs in AG-ReID.v1 [20] datasets is far lower than in-ground ReID datasets like Market1501 [40] and DukeMTMC-reID [41], highlighting the lack of data resources in this field. Although AG-ReID.v2 [21] attempts to address this shortfall by expanding the dataset, the large number of wearable device images included in the new data does not substantially enhance coverage of the differences between aerial and ground views, affecting its practical value. Additionally, synthetic datasets like CARGO [36], while approaching real datasets in scale, lack real-world complexity, limiting model performance in actual scenarios. Therefore, developing the larger-scale, more representative aerial-ground cross-view ReID dataset, and designing efficient algorithms to utilize these datasets effectively, become urgent problems to be solved in this field. To overcome this problem, we contribute two real-world large-scale AGPReID datasets: LAGPeR and G2APS-ReID.
@@ 4.1 LAGPeR
Data Division and Experimental Setup: For the LAGPeR dataset, we select 12 cameras (including 8 ground cameras and 4 drone cameras) from the first four scenes as the training set, while images from 9 cameras in the remaining three scenes are used for evaluation. To simulate real-world disturbances, we include IDs that appeare only in one camera as noise items in the dataset. After filtering, we identify $1,523$ IDs for evaluation, while the remaining $2,708$ IDs are used as the training set. During the data prepossessing stage, we aim to select the most representative images for each ID as query images. To achieve this, we employ a method based on gradient histograms and K-nearest neighbor clustering algorithms. Specifically, we calculate the gradient histogram features of images with the same ID and view and use K-nearest neighbor clustering to divide the images into K groups, randomly selecting one image from each group as a query image, thus selecting K representative images as queries. Ultimately, we successfully select $3,046$ images of $1,523$ IDs from each perspective as query images, constructing a comprehensive task setup (as shown in Tab. 2). In terms of the experimental setup, compared to conventional AGPReID datasets, we add $G\rightarrow A+G$ setting, which includes images from both ground and aerial viewpoints in the gallery to more comprehensively evaluate the model’s performance.
@@ 6 Conclusion
This paper focuses on cross-view ReID, specifically AGPReID. Firstly, we propose the Self-Calibrating and Adaptive Prompts (SeCap) method to address the significant view differences in the AGPReID task. By re-calibrate prompts to match the current view based on the input adaptively, the SeCap significantly enhances model performance. Secondly, we contribute two real-world large-scale AGPReID datasets, LAGPeR and G2APS-ReID. The former is collected and annotated by us independently, covering $4,231$ unique identities and containing $63,841$ images; the latter is reconstructed from the person search dataset G2APS. Finally, the experiments on AGPReID datasets demonstrate the superiority of our method.

### v2

@@ 1 Introduction
Person re-identification (ReID), as the cornerstone of intelligent surveillance systems, fundamentally relies on accurately identifying individuals across different camera viewpoints and complex environmental changes [11, 12, 25, 37, 23, 29]. However, traditional ReID methods are often limited to the same view, such as ground view [34, 41, 27, 35, 36, 24] or aerial view [39, 14, 4, 10, 28], and fail to adequately address the challenges posed by extreme visual transformations and the integration of complementary information in cross-view scenarios (e.g., combining ground and aerial cameras). These challenges are more prevalent in real-world applications, thereby introducing the problem of the Aerial-Ground Person Re-Identification (AGPReID).
@@ 2.3 Aerial-Ground ReID Datasets
In the AGPReID field, the scarcity and limitations of datasets become the key factors constraining research progress. Compared to mature view-invariant datasets [42, 43], AGPReID datasets are scarce in number and far from meeting research needs in terms of real-world data scale. For instance, the number of identity IDs in AG-ReID.v1 [20] datasets is far lower than in-ground ReID datasets like Market1501 [42] and DukeMTMC-reID [43], highlighting the lack of data resources in this field. Although AG-ReID.v2 [21] attempts to address this shortfall by expanding the dataset, the large number of wearable device images included in the new data does not substantially enhance coverage of the differences between aerial and ground views, affecting its practical value. Additionally, synthetic datasets like CARGO [38], while approaching real datasets in scale, lack real-world complexity, limiting model performance in actual scenarios. Therefore, developing the larger-scale aerial-ground cross-view ReID dataset, and designing efficient algorithms to utilize these datasets effectively, become urgent problems to be solved in this field. To overcome this problem, we contribute two real-world large-scale AGPReID datasets: LAGPeR and G2APS-ReID.
@@ 4.1 LAGPeR
Data Division and Experimental Setup: For the LAGPeR dataset, we select 12 cameras (including 8 ground cameras and 4 drone cameras) from the first four scenes as the training set, while images from 9 cameras in the remaining three scenes are used for evaluation. To simulate real-world disturbances, we include IDs that appeare only in one camera as noise items in the dataset. After filtering, we identify $1,523$ IDs for evaluation, while the remaining $2,708$ IDs are used as the training set. During the data prepossessing stage, we aim to select the most representative images for each ID as query images. To achieve this, we employ a method based on gradient histograms and K-nearest neighbor clustering algorithms. Ultimately, we successfully select $3,046$ images of $1,523$ IDs from each perspective as query images, constructing a comprehensive task setup (as shown in Tab. 2). In terms of the experimental setup, compared to conventional AGPReID datasets, we add $G\rightarrow A+G$ setting, which includes images from both ground and aerial viewpoints in the gallery to evaluate the model’s performance more comprehensively.
@@ 6 Conclusion
This paper focuses on cross-view ReID, specifically AGPReID. Firstly, we propose the Self-Calibrating and Adaptive Prompts (SeCap) method to address the significant view differences. By re-calibrate prompts to match the current view based on the input adaptively, the SeCap significantly enhances model performance. Secondly, we contribute two real-world large-scale AGPReID datasets, LAGPeR and G2APS-ReID. Finally, the experiments on AGPReID datasets demonstrate the superiority of our method.
Acknowledgments.This work is supported by the National Natural Science Foundation of China (U23B2013) and the Guangdong Basic and Applied Basic Research Foundation (2025A1515011465).
@@ 13 Broader impact
14 Scene examples of the LAGPeR dataset.
As Fig.9, we show some scene examples of the LAGPeR dataset, where A-Cam represents the aerial view camera and G-Cam represents the ground view camera. Each column represents the same scene.
Figure 9: Example images from the LAGPeR, where each column represents images from the same scene.
15 Performance of $G\leftrightarrow G$
[…]
Method ViT VDT TransReID SeCap
Rank-1 78.36 77.18 79.23 79.57
mAP 77.38 76.35 78.09 78.26